### v1

@@ 6 Real data analysis
In this section, we analyze the AUCS dataset (Rossi and Magnani, 2015), which is a public dataset collected by the employees at the Department of Computer Science of Aarhus University, to demonstrate the utility of the proposed algorithm. Though privacy preservation is prominent nowadays, to the best of our knowledge, networks known to be noisy and perturbed by the RR mechanism are not currently available to the public. Hence, we regard the AUCS dataset as the original data without any perturbation. We will mimic the actual setup to see how the privacy and local computation affects the clustering performance of the proposed algorithm.
@@
There are 61 employees (out of the total number of 142) who decided to join the survey, consisting of professors, postdoctoral researchers, Ph.D. students, and administration staff. Five types of relationships among the employees are measured. Namely, current working relationships, repeated leisure activities, regularly eating lunch together, co-authorship of a publication, and friendship on Facebook. The first three relationships were collected via a questionnaire. All the respondents answered all questions indicating that the data was complete. And the last two relationships were acquired online via the DBLP database and Facebook. As a result, the original dataset consists of 5 undirected networks with 61 common nodes. Besides the relationships, the research group labels that 55 out of 61 employees are available. There are eight research groups ($K=8$) in total. Our analysis uses the subnetworks of 55 employees with known labels.
@@
The topology of networks with respect to different relationships is shown in Figure 2 with colors indicating the research groups. It is clear to see that the network structures are quite different. To take a closer look at the community structures, we reorder the nodes according to their underlying research groups (i.e., nodes corresponding to the same research group are aligned next to each other). We further plot the corresponding network adjacency matrices in Figure 3, where the whole network (Figure 3(a)) corresponds to the network that contains the edges of all five relationships. The whole network shows conspicuous block structures in that nodes in the same research group have more links than those across different research groups. Hence, it is reliable to use the research group labels as the underlying community memberships. Distinct from the whole network, we see that the community structures of each network layer are heterogeneous in that some communities are not evident or even missing to different extents.
(a) Work
(b) Facebook
(c) Leisure
(d) Lunch
(e) Coauthor
Figure 2: A visualization of the AUCS network where vertex colors represent research groups. Each network corresponds to a specific relationship among the vertices.
@@
  • Effect of $L$: $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 5.
  • Effect of $q$: $L=5$, $q^{\prime}=0.95$, and $q$ ranges from 0.6 to 1.
  • Effect of $q^{\prime}$: $L=5$, $q=0.95$, and $q^{\prime}$ ranges from 0.6 to 1.
@@ Results.
The averaged misclassification rates over 20 replications are displayed in Figure 4. We have the following observations. First, as $L$ increases, the misclassification rate of all four methods decreases, showing the benefits of aggregating the information of local networks. Note that when $L=5$, a small drop in terms of the clustering performance appears, partially because the fifth network (i.e., the coauthor network) is rather sparse and has a weak signal of communities. Second, as expected, as $q$ and $q^{\prime}$ increase, the clustering performance of all four methods becomes better. We note that the clustering performance with $q^{\prime}$ fixed to be 0.95 (See Figure 4(c)) is better than that with $q$ fixed to be 0.95 (See Figure 4(d)), which shows that large $q^{\prime}$ is more desirable than large $q$. In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0. This is consistent with the theoretical results. Third, consistent with the simulation results, the proposed method ppDSC is better than the biased method ppDSC-1b and ppDSC-2b and is close to the baseline method Oracle as $q$ and $q^{\prime}$ increase.
(a) Whole
(b) Work
(c) Facebook
(d) Leisure
(e) Lunch
(f) Coauthor
Figure 3: A visualization of the network adjacency matrices of the AUCS dataset. The nodes are permutated according to their underlying research groups. (a) corresponds to the network adjacency matrix that contains all the edges of the five relationships. (b)-(f) correspond to the adjacency matrices of the five relationships among vertices, respectively.
@@
Figure 4: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the AUCS network. The effect of the privacy parameters $q,q^{\prime}$ ($q,q^{\prime}$ synchronously vary, $q^{\prime}$ fixed but $q$ varies, $q$ fixed but $q^{\prime}$ varies), and the number of networks $L$ are shown, respectively. The misclassification rate with respect to the research group labels are evaluated.
@@ 7 Conclusion
Driven by practical data collection and the associated analytic challenges, we provide a powerful toolbox for community detection and estimation in multi-layer SBMs based on locally stored networks with privacy protection. In the toolbox, a new algorithm named ppDSC has been developed, and it is shown to be both communication-efficient and privacy-aware. We also propose a two-step bias-adjustment procedure to remove the biases caused by the RR perturbation and take the square of the private-preserved network matrices. The procedure is found to be effective both theoretically and numerically. Moreover, we have rigorously investigated the misclassification rate of the algorithm. We show how the privacy parameters affect the error bound. In particular, we demonstrate the effect of heterogeneity among population networks and the asymmetry roles of the privacy parameters $q$ and $q^{\prime}$ on the estimation accuracy through rigorous theoretical justifications. The numerical experiments well support our theoretical findings.
The proposed methodology can be extended to a variety of scenarios. First, the ppDSC algorithm involves multiple steps. The methods used in some steps are quite general and can be applied in other contexts. For example, the proposed bias-adjustment procedure can be used in other stochastic block models in different distributed SVD frameworks, e.g., Fan et al. (2019); Arroyo et al. (2021). Second, it is well-known that privacy could be amplified by sampling (Balle et al., 2018). With the same level of noise from privacy preservation, a DP algorithm with sampled data leaks less privacy than that with all data. Therefore, it is desirable to study the trade-off between privacy gain and information loss when sampling the network entries in network analysis. Third, for statistical convenience, we consider the local-DP (Duchi et al., 2018) protection of networks. We may extend the proposed methods in the context of central DP. Fourth, we focus on the distributed learning framework with only one round of communication. It is desirable to extend the results to the distributed framework with multiple rounds, say the communication-efficient distributed power method (Li et al., 2021). Last but not least, the method and theory for estimating $K$ can be an interesting topic to explore in the future. (Fishkind et al., 2013; Li et al., 2020; Ma et al., 2021).
@@
Appendix A provides the details of the bias-adjustment procedure. Appendix B contains the additional theoretical results and discussions. Appendix C contains the additional simulation experiment. Appendix D contains the technical proofs for Theorems 1-4 and Corollary 1. Several auxiliary lemmas are provided in Appendix E.
@@ Appendix A Details of bias-adjustment
Appendix B Additional theoretical results and discussions
B.1 Bound for $\overline{M}_{l}$
We provide the bound for the single matrix $\overline{M}_{l}$. As we will see, due to the absence of $L$, the condition and the dominating term in the resulting bound compared with the centralized bound is changed. The results are summarized in Theorem B.1.
Theorem B.1
Suppose Assumption 2 holds, and $\epsilon\lesssim{\log}\,n$, the estimator $\overline{M}_{l}$ for $Q_{l}$ satisfies
$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$ (B.1)
with probability larger than $1-O(Ln^{-\gamma})$, where $\gamma,C>0$ are some fixed constants.
The proof of Theorem B.1 is included in Appendix D.
B.2 The effect of bias-adjustment
@@
$\scriptsize\|\tilde{M}-\tilde{Q}\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}$ (B.2)
@@
The proof is similar to that of Theorem 3 provided in Appendix D; hence we omit it. It can be easily seen that $\mathcal{I}$ is larger than (3) in Theorem 3. Hence, if $q,q^{\prime}>c$ for some $c>1/2$, which implies that the multiplier $(q+q^{\prime}-1)^{2}$ is a constant, then (B.2) and (B.3) are both inferior than (3). To derive the bound for the eigen-vector and the misclassification error bound, we note that we can either use (B.2) or (B.3). What we need is only the minimum non-zero eigen-value of the corresponding population matrix. Specifically, if we use (B.3), then the population matrix is $Q$, which is the same as our setting. While if we use (B.2), then the minimum non-zero eigen-value of $\tilde{Q}$ is required. Here, we consider the two-parameter SBMs that $P_{l}$’s are homogeneous, with the inter-community and cross-community probabilities being $a$ and $b$, respectively. In this regime, the minimum non-zero eigen-value of $(q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}$ is identical to that of $(q+q^{\prime}-1)P_{l}$, which implies that the error for the eigen-vector (counterpart of Corollary 1) and the misclassification error (counterpart of Theorem 4) totally depend on (B.3). As a result, we can also see the accuracy gain in eigen-vector estimation and clustering.
@@ Appendix C Additional numerical analysis and results
In Theorem 2, we show that the eigen-gap (i.e., $\lambda_{{\min}}(Q)$) divided by the heterogeneity of connectivity matrices (i.e., $\mathcal{H}$ in Assumption 4) can affect the estimation accuracy of the proposed algorithm theoretically. The heterogeneity can lead to either accuracy gain or loss, depending on the model structure of the network. To illustrate this phenomenon numerically, we consider the following two models for demonstrating the curses of heterogeneity and the blessings of heterogeneity, respectively.
  • Model I (Curses of heterogeneity). We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. The networks are generated from multi-layer SBMs and equally belong to the following four different classes,
@@
    For this model, one can imagine that larger $\alpha$ indicates larger heterogeneity but a similar level of signal strength (eigen-gap). Therefore, larger $\alpha$ leads to larger scaled heterogeneity (i.e., eigen-gap divided by heterogeneity; see the discussions after Theorem 2 for the details). This is verified in Figure C.1(a).
  • Model II (Blessings of heterogeneity). We fix $K=2$, $L=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. Similar to Arroyo et al. (2021), the networks are generated from SBMs and equally belong to the following four different classes,
@@
    For this model, larger $\alpha$ indicates larger heterogeneity but stronger signal strength (eigen-gap). Thus, the composite effects lead to larger $\alpha$ corresponding to smaller scaled heterogeneity. See Figure C.2(a) for the illustration.
@@ Results.
The averaged results over 20 replications for Model I and II are displayed in Figure C.1 and C.2, respectively. The results are consistent with the theoretical results given in Theorem 2. For Model I, we see that large heterogeneity (i.e., large $\alpha$) leads to inferior performance of ppDSC, especially in terms of the projection distance, which is partially because that large $\alpha$ leads to large scaled heterogeneity (i.e., eigen-gap divided by heterogeneity) in Theorem 2; see Figure C.1. For Model II, the results are on the opposite side in that large heterogeneity (i.e., large $\alpha$) leads to performance gain of ppDSC, which is partially due to the fact that large $\alpha$ leads to large eigen-gap and small scaled heterogeneity; see Figure C.2. In addition, our estimator ppDSC is better than the biased ppDSC-1b and ppDSC-2b, and is closer to the Oracle estimator.
@@
Figure C.1: Illustration for the curse of heterogeneity (Model I). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.
@@
Figure C.2: Illustration for the blessing of heterogeneity (Model II). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b and Oracle.

### v2

@@ 6 Real data analysis
In this section, we analyze the AUCS dataset (Rossi and Magnani, 2015), which is a public dataset collected by the employees at the Department of Computer Science of Aarhus University, to demonstrate the utility of the proposed algorithm. Though privacy preservation is prominent nowadays, to the best of our knowledge, networks known to be noisy and perturbed by the RR mechanism are not currently available to the public. Hence, we regard the AUCS dataset as the original data without any perturbation. We will mimic the actual setup to see how the privacy and local computation affect the clustering performance of the proposed algorithm.
@@
There are 61 employees (out of the total number of 142) who decided to join the survey, consisting of professors, postdoctoral researchers, Ph.D. students, and administration staff. Five types of relationships among the employees are measured. Namely, current working relationships, repeated leisure activities, regularly eating lunch together, co-authorship of a publication, and friendship on Facebook. The first three relationships were collected via a questionnaire. All the respondents answered all questions, indicating that the data was complete. And the last two relationships were acquired online via the DBLP database and Facebook. As a result, the original dataset consists of 5 undirected networks with 61 common nodes. Besides the relationships, the research group labels that 55 out of 61 employees are available. There are eight research groups ($K=8$) in total. Our analysis uses the subnetworks of 55 employees with known labels.
@@
The topology of networks with respect to different relationships is shown in Figure C.4 in the Supplementary Materials with colors indicating the research groups. It is clear to see that the network structures are quite different. To take a closer look at the community structures, we reorder the nodes according to their underlying research groups (i.e., nodes corresponding to the same research group are aligned next to each other). We further plot the corresponding network adjacency matrices (see Figure C.5 in the Supplementary Materials), where the whole network (Figure C.5(a)) corresponds to the network that contains the edges of all five relationships. The whole network shows conspicuous block structures in that nodes in the same research group have more links than those across different research groups. Hence, it is reliable to use the research group labels as the underlying community memberships. Distinct from the whole network, we see that the community structures of each network layer are heterogeneous in that some communities are not evident or even missing to different extents.
@@
Consistent with the simulation experiments, we compare our method ppDSC with ppDSC-1b, ppDSC-2b, ppSC and Oracle. The detailed description of each method can be found in Section 5. We study how the performance of each method alters with the network number $L$ and the privacy parameters $q$ and $q^{\prime}$. Specifically, we fix $q(q^{\prime})$ to be 0.95 and vary $q^{\prime}(q)$ to verify the different effect of $q$ and $q^{\prime}$ indicated by the theory. To save space, the detailed descriptions on the experimental set-up is relegated to Section C of the Supplementary Materials.
@@ Results.
The averaged misclassification rates over 20 replications are displayed in Figure 4. We have the following observations. First, as $L$ increases, the misclassification rate of all five methods decreases, showing the benefits of aggregating the information of local networks. Note that when $L=5$, a small drop in terms of the clustering performance appears, partially because the fifth network (i.e., the coauthor network) is rather sparse and has a weak signal of communities. Second, as expected, as $q$ and $q^{\prime}$ increase, the clustering performance of all four privacy-preserving methods becomes better. We note that the clustering performance with $q^{\prime}$ fixed to be 0.95 (See Figure 4(c)) is better than that with $q$ fixed to be 0.95 (See Figure 4(d)), which shows that large $q^{\prime}$ is more desirable than large $q$. In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0. This is consistent with the theoretical results. Third, consistent with the simulation results, the proposed method ppDSC is better than the biased method ppDSC-1b and ppDSC-2b and is close to the baseline methods ppSC and Oracle as $q$ and $q^{\prime}$ increase.
@@
Figure 4: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the AUCS network. The effect of the privacy parameters $q,q^{\prime}$ ($q,q^{\prime}$ synchronously vary, $q^{\prime}$ fixed but $q$ varies, $q$ fixed but $q^{\prime}$ varies), and the number of networks $L$ are shown, respectively. The misclassification rate with respect to the research group labels is evaluated.
@@ 7 Conclusion
Driven by practical data collection and the associated analytic challenges, we provide a powerful toolbox for community detection and estimation in multi-layer SBMs based on locally stored networks with privacy protection. In the toolbox, a new algorithm named ppDSC has been developed, and it is shown to be both communication-efficient and privacy-aware. We also propose a two-step bias-adjustment procedure to remove the biases caused by the RR perturbation and taking the square of the debiased adjacency matrices. The procedure is found to be effective both theoretically and numerically. Moreover, we have rigorously investigated the misclassification rate of the algorithm. In particular, we show that when the number of machines and the effective heterogeneity are well-controlled, the misclassification rate of our distributed algorithm matches its centralized counterpart. More interestingly, our result indicates the best choice of privacy parameters under a given privacy budget $\epsilon$. In addition, we demonstrate the effect of heterogeneity among population networks and the asymmetrical roles of the privacy parameters $q$ and $q^{\prime}$ on the estimation accuracy. The numerical experiments support our theoretical findings well.
The proposed methodology can be extended to a variety of scenarios. First, the ppDSC algorithm involves multiple steps. The methods used in some steps are quite general and can be applied in other contexts. For example, the proposed bias-adjustment procedure can be used in other stochastic block models in different distributed SVD frameworks, e.g., Fan et al. (2019); Arroyo et al. (2021). The method for computing the local eigenspaces can also be replaced with the tensor-based method, e.g., Jing et al. (2021). Second, it is well-known that privacy could be amplified by sampling (Balle et al., 2018). With the same level of noise from privacy preservation, a DP algorithm with sampled data leaks less privacy than that with all data. Therefore, it is desirable to study the trade-off between privacy gain and information loss when sampling the network entries in network analysis. Third, for statistical convenience, we consider the local-DP (Duchi et al., 2018) protection of networks. We may extend the proposed methods in the context of central DP. Fourth, we focus on the distributed learning framework with only one round of communication. It is desirable to extend the results to the distributed framework with multiple rounds, say the communication-efficient distributed power method (Guo et al., 2024). Last but not least, the method and theory for estimating $K$ can be an interesting topic to explore in the future. (Fishkind et al., 2013; Li et al., 2020; Ma et al., 2021).
@@
Section A provides the details of the bias-adjustment procedure. Section B includes the additional theoretical results and discussions. Section C presents the additional simulation experiments and the results of real data analysis. Section D contains the technical proofs. Several auxiliary lemmas are provided in Section E.
@@ Appendix A Details of bias-adjustment
Appendix B The effect of bias-adjustment
@@
$\|\tilde{M}-\tilde{Q}\|_{2}\lesssim(q+q^{\prime}-1)^{2}\mathcal{I}$ (B.1)
@@
The proof is similar to that of Theorem 3 provided in Appendix D; hence we omit it. It can be easily seen that $\mathcal{I}$ is larger than (3) in Theorem 3. Hence, if $q,q^{\prime}>c$ for some $c>1/2$, which implies that the multiplier $(q+q^{\prime}-1)^{2}$ is a constant, then (B.1) and (B.2) are both inferior than (3). To derive the bound for the eigen-vector and the misclassification error bound, we note that we can either use (B.1) or (B.2). What we need is only the minimum non-zero eigen-value of the corresponding population matrix. Specifically, if we use (B.2), then the population matrix is $Q$, which is the same as our setting. While if we use (B.1), then the minimum non-zero eigen-value of $\tilde{Q}$ is required. Here, we consider the two-parameter SBMs that $P_{l}$’s are homogeneous, with the inter-community and cross-community probabilities being $a$ and $b$, respectively. In this regime, the minimum non-zero eigen-value of $(q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}$ is identical to that of $(q+q^{\prime}-1)P_{l}$, which implies that the error for the eigen-vector (counterpart of Theorem 4) and the misclassification error (counterpart of Theorem 5) totally depend on (B.2). As a result, we can also see the accuracy gain in eigen-vector estimation and clustering.
@@ C.1 Effect of heterogeneity
In Theorem 2, we show that the eigen-gap (i.e., $\lambda_{{\min}}(Q)$) divided by the heterogeneity of connectivity matrices (i.e., $\mathcal{H}$ in Definition 2) can affect the estimation accuracy of the proposed algorithm theoretically. The heterogeneity can lead to either accuracy gain or loss, depending on the model structure of the network. To illustrate this phenomenon numerically, we consider the following two models for demonstrating the effect of heterogeneity.
  • Model I. We fix $K=2$, $L=m=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. The networks are generated from multi-layer SBMs and equally belong to the following four different classes,
@@
    For this model, one can imagine that larger $\alpha$ indicates larger heterogeneity but a similar level of signal strength (eigen-gap). Therefore, larger $\alpha$ leads to larger effective heterogeneity (i.e., eigen-gap divided by heterogeneity; see the discussions after Theorem 2 for the details). This is verified in Figure C.1(a).
  • Model II. We fix $K=2$, $L=m=12$, $q,q^{\prime}=0.8$ and let $n$ vary from 20 to 400. The communities are balanced. Similar to Arroyo et al. (2021), the networks are generated from SBMs and equally belong to the following four different classes,
@@
    For this model, larger $\alpha$ indicates larger heterogeneity but stronger signal strength (eigen-gap). Thus, the composite effects lead to larger $\alpha$ corresponding to smaller effective heterogeneity. See Figure C.2(a) for the illustration.
@@ Results.
The averaged results over 20 replications for Model I and II are displayed in Figure C.1 and C.2, respectively. The results are consistent with the theoretical results given in Theorem 2. For Model I, we see that large heterogeneity (i.e., large $\alpha$) leads to inferior performance of ppDSC, especially in terms of the projection distance, which is partially because large $\alpha$ leads to large effective heterogeneity (i.e., eigen-gap divided by heterogeneity) in Theorem 2; see Figure C.1. For Model II, the results are on the opposite side in that large heterogeneity (i.e., large $\alpha$) leads to performance gain of ppDSC, which is partially due to the fact that large $\alpha$ leads to large eigen-gap and small effective heterogeneity; see Figure C.2. In addition, our estimator ppDSC is better than the biased ppDSC-1b and ppDSC-2b, and is closer to the baseline estimators ppSC and Oracle estimator.
@@
Figure C.1: Illustration for the negative effect of heterogeneity (Model I). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle.
@@
Figure C.2: Illustration for the positive effect of heterogeneity (Model II). (a) corresponds to population parameters including the heterogeneity, eigen-gap and their ratios against $\alpha$. (b) and (c) correspond to the projection distance and misclassification rate of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle.
C.2 Effect of network sparsity
We evaluate the effect of network sparsity on the performance of ppDSC under different privacy budgets. To incorporate the network sparsity, we generate $B_{l}$’s by $B_{l}:=0.8\gamma B^{(1)}$ for $l=1,...,\lfloor L/2\rfloor$ and $B_{l}:=0.6\gamma B^{(2)}$ for $l=\lfloor L/2\rfloor+1,...,L$ with varying $\gamma$. Other parameters are $n=210$ and $L=m=12$. The averaged results over 20 replications under different $q$’s are displayed in Figure C.3. We assume $q,q^{\prime}$ relates to $\epsilon$ via $q=q^{\prime}=\frac{{\rm e}^{\epsilon}}{1+{\rm e}^{\epsilon}}$. Under a given privacy budget, a denser network yields better clustering performance. In other words, under a given network sparsity, a denser network permits a wider range of privacy budget to attain a specified misclassification rate or projection distance.
(a) Projection distance
(b) Misclassification rate
Figure C.3: The performance of ppDSC with varying sparsity under different $q,q^{\prime}$’s.
C.3 Experimental details and additional figures of real data analysis
Figures C.4 and C.5 show the visualization of the AUCS dataset in terms of networks and matrices, respectively.
(a) Work
(b) Facebook
(c) Leisure
(d) Lunch
(e) Coauthor
Figure C.4: A visualization of the AUCS network where vertex colors represent research groups. Each network corresponds to a specific relationship among the vertices.
(a) Whole
(b) Work
(c) Facebook
(d) Leisure
(e) Lunch
(f) Coauthor
Figure C.5: A visualization of the network adjacency matrices of the AUCS dataset. The nodes are permutated according to their underlying research groups. (a) corresponds to the network adjacency matrix that contains all the edges of the five relationships. (b)-(f) correspond to the adjacency matrices of the five relationships among vertices, respectively.
We next provide the details of the methods compared in Section 6.
Methods compared in Section 6.
  • Effect of $L$: $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 5.
  • Effect of $q$: $L=5$, $q^{\prime}=0.95$, and $q$ ranges from 0.6 to 1.
  • Effect of $q^{\prime}$: $L=5$, $q=0.95$, and $q^{\prime}$ ranges from 0.6 to 1.